# Date of Acceptance: Nov. 9th 2025

### v1

@@ 6 Conclusion
Lemma A.1 (Stability of Dominant Subspace; see Theorem 1 in [o2023matrices]).
@@
Let $p$ and $p^{\prime}$ be the normalized singular value spectra of $X$ and $X^{\prime}$. This redistribution of squared singular values, where the total sum increases while the dominant part remains roughly constant, leads to a flattening of the normalized singular value distribution $p^{\prime}$ compared to $p$. Specifically, the increase in total mass ($\|X^{\prime}\|_{F}^{2}>\|X\|_{F}^{2}$) is concentrated in the tail, which reduces the probability mass of the dominant components. Since the Shannon entropy $H(p)$ is Schur-concave, it increases when the probability mass is transferred from larger components to smaller ones [marshall1979inequalities]. Therefore:
@@ C.2 Action anticipation Module Architecture
Future queries as learnable representations. Instead of relying on a single deterministic output for action anticipation, we introduce learnable queries—denoted as future queries ($\mathbf{Q}_{\text{future}}$) to predict multiple action possibilities [gong2022future]. These queries are randomly initialized and learned during training. Given $\mathbf{N_{q}}$ future queries, we define: $\mathbf{Q}_{\text{future}}\in\mathbb{R}^{N_{q}\times C}$ where $C$ is the hidden dimension. These queries are independent of any specific frame and instead act as a learnable representation that extracts relevant information from the past observations.
@@ D.1 Datasets
We aim to evaluate R3D’s ability to anticipate human actions in diverse and realistic scenarios, particularly in settings where spatial reasoning and multi-modal fusion play a critical role. To this end, we utilize three action anticipation datasets: NTURGBD [shahroudy2016ntu], UTKinect-Action3D [xia2012view], and DARai [kaviani2025hierarchical]. The UTKinect-Action3D dataset was collected by the University of Texas at Austin and features 10 types of human actions, each performed twice by 10 subjects from multiple viewpoints. The dataset poses additional challenges due to actor-dependent variations, occlusions caused by human-object interactions, and body parts moving out of the field of view. The NTURGBD dataset is one of the largest and most diverse action recognition datasets, designed for 3D human action analysis. It contains 56,880 video samples covering 60 action classes performed by 40 subjects from 80 viewpoints. The DARai dataset offers highly realistic scenarios that closely resemble real-world human behavior. Unlike UTKinect and NTURGBD, the videos in DARai are untrimmed, showcasing raw, continuous human activity in real-world contexts without artificial segmentation. It comprises 150 action classes, 50 participants, two distinct exocentric views, and three levels of hierarchical labels.
@@ D.2 Experimental setups
We use pretrained ResNet features as input visual features for the NTURGBD, UTKinects, and DARai. To align with the temporal resolution of each dataset, the sampling rates are set to 15 for DARai, and 1 for UTKinects and NTURGBD. We use the number of Future Queries $\mathbf{N_{\text{q}}}$ fixed at 8. Based on the density of each dataset, the hidden dimension size $D$ is set to 128 for both NTURGBD, UTKinects, and DARai. During training, the observation rate $\alpha$ is set to $\alpha\in\{0.2,0.3,0.5\}$, while the prediction rate $\beta$ is fixed at 0.5. Yet when training UTKinects, to further augment the training dataset, we use the observation rate $\alpha\in\{0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55\}$. The model is trained for 60 epochs using the AdamW optimizer [loshchilov2017decoupled] with a learning rate of 1e31e-3 and a batch size of 8. A cosine annealing warm-up scheduler [loshchilov2016sgdr] is applied, with warm-up stages spanning 10 epochs. For evaluation setup, we follow the long-term action anticipation framework protocol [abu2018will, abu2021long, sener2020temporal, ke2019time]. We set the observation rate $\alpha\in\{0.2,0.3\}$ and prediction rate $\beta\in\{0.1,0.2,0.3,0.5\}$, and measure mean over classes (MoC) accuracy for evaluation metrics. To ensure consistency, we report average performance across 3 number of iteration, each with fixed seeds $1,10,13452$.
@@ D.3 Computational cost and Scalabilty
GTAN [zatsarynna2024gated] 5.920 49 GFLOPs
@@
Computational cost. We evaluate the computational cost and scalability of R3D on an NVIDIA A40 GPU. For comparison, we chose GTAN [zatsarynna2024gated], the current state-of-the-art model reported in Table 1. As shown in Table 7, R3D processes a single RGB-Depth frame in 0.119 ms, requiring 61.77MB of memory and 0.58 GFLOPs. In contrast, GTAN requires 5.92ms per RGB frame and approximately 49 GFLOPs. This gap arises because GTAN is a diffusion-based model whose computational cost scales linearly with the number of sampling steps, while R3D avoids such overhead.

### v2

@@ 6 Conclusion
Lemma A.1 (Stability of Dominant Subspace; see Theorem 1 in [52]).
@@
Let $p$ and $p^{\prime}$ be the normalized singular value spectra of $X$ and $X^{\prime}$. This redistribution of squared singular values, where the total sum increases while the dominant part remains roughly constant, leads to a flattening of the normalized singular value distribution $p^{\prime}$ compared to $p$. Specifically, the increase in total mass ($\|X^{\prime}\|_{F}^{2}>\|X\|_{F}^{2}$) is concentrated in the tail, which reduces the probability mass of the dominant components. Since the Shannon entropy $H(p)$ is Schur-concave, it increases when the probability mass is transferred from larger components to smaller ones [51]. Therefore:
@@ C.2 Action anticipation Module Architecture
Future queries as learnable representations. Instead of relying on a single deterministic output for action anticipation, we introduce learnable queries—denoted as future queries ($\mathbf{Q}_{\text{future}}$) to predict multiple action possibilities [22]. These queries are randomly initialized and learned during training. Given $\mathbf{N_{q}}$ future queries, we define: $\mathbf{Q}_{\text{future}}\in\mathbb{R}^{N_{q}\times C}$ where $C$ is the hidden dimension. These queries are independent of any specific frame and instead act as a learnable representation that extracts relevant information from the past observations.
@@ D.1 Datasets
We aim to evaluate R3D’s ability to anticipate human actions in diverse and realistic scenarios, particularly in settings where spatial reasoning and multi-modal fusion play a critical role. To this end, we utilize three action anticipation datasets: NTURGBD [59], UTKinect-Action3D [65], and DARai [32]. The UTKinect-Action3D dataset was collected by the University of Texas at Austin and features 10 types of human actions, each performed twice by 10 subjects from multiple viewpoints. The dataset poses additional challenges due to actor-dependent variations, occlusions caused by human-object interactions, and body parts moving out of the field of view. The NTURGBD dataset is one of the largest and most diverse action recognition datasets, designed for 3D human action analysis. It contains 56,880 video samples covering 60 action classes performed by 40 subjects from 80 viewpoints. The DARai dataset offers highly realistic scenarios that closely resemble real-world human behavior. Unlike UTKinect and NTURGBD, the videos in DARai are untrimmed, showcasing raw, continuous human activity in real-world contexts without artificial segmentation. It comprises 150 action classes, 50 participants, two distinct exocentric views, and three levels of hierarchical labels.
@@ D.2 Experimental setups
We use pretrained ResNet features as input visual features for the NTURGBD, UTKinects, and DARai. To align with the temporal resolution of each dataset, the sampling rates are set to 15 for DARai, and 1 for UTKinects and NTURGBD. We use the number of Future Queries $\mathbf{N_{\text{q}}}$ fixed at 8. Based on the density of each dataset, the hidden dimension size $D$ is set to 128 for both NTURGBD, UTKinects, and DARai. During training, the observation rate $\alpha$ is set to $\alpha\in\{0.2,0.3,0.5\}$, while the prediction rate $\beta$ is fixed at 0.5. Yet when training UTKinects, to further augment the training dataset, we use the observation rate $\alpha\in\{0.2,0.25,0.3,0.35,0.4,0.45,0.5,0.55\}$. The model is trained for 60 epochs using the AdamW optimizer [49] with a learning rate of 1e31e-3 and a batch size of 8. A cosine annealing warm-up scheduler [50] is applied, with warm-up stages spanning 10 epochs. For evaluation setup, we follow the long-term action anticipation framework protocol [1, 2, 58, 34]. We set the observation rate $\alpha\in\{0.2,0.3\}$ and prediction rate $\beta\in\{0.1,0.2,0.3,0.5\}$, and measure mean over classes (MoC) accuracy for evaluation metrics. To ensure consistency, we report average performance across 3 number of iteration, each with fixed seeds $1,10,13452$.
@@ D.3 Computational cost and Scalabilty
GTAN [72] 5.920 49 GFLOPs
@@
Computational cost. We evaluate the computational cost and scalability of R3D on an NVIDIA A40 GPU. For comparison, we chose GTAN [72], the current state-of-the-art model reported in Table 1. As shown in Table 7, R3D processes a single RGB-Depth frame in 0.119 ms, requiring 61.77MB of memory and 0.58 GFLOPs. In contrast, GTAN requires 5.92ms per RGB frame and approximately 49 GFLOPs. This gap arises because GTAN is a diffusion-based model whose computational cost scales linearly with the number of sampling steps, while R3D avoids such overhead.